# MM-Soc: Benchmarking Multimodal Large Language Models in Social Media Platforms

### v1

@@ Abstract
Social media platforms are hubs for multimodal information exchange, encompassing text, images, and videos, making it challenging for machines to comprehend the information or emotions associated with interactions in online spaces. Multimodal Large Language Models (MLLMs) have emerged as a promising solution to address these challenges, yet struggle with accurately interpreting human emotions and complex contents like misinformation. This paper introduces MM-Soc, a comprehensive benchmark designed to evaluate MLLMs’ understanding of multimodal social media content. MM-Soc compiles prominent multimodal datasets and incorporates a novel large-scale YouTube tagging dataset, targeting a range of tasks from misinformation detection, hate speech detection, and social context generation. Through our exhaustive evaluation on ten size-variants of four open-source MLLMs, we have identified significant performance disparities, highlighting the need for advancements in models’ social understanding capabilities. Our analysis reveals that, in a zero-shot setting, various types of MLLMs generally exhibit difficulties in handling social media tasks. However, MLLMs demonstrate performance improvements post fine-tuning, suggesting potential pathways for improvement.
@@ 1 Introduction
Social media platforms have become the epicenter of multimodal information exchange, blending various formats of content such as text, images, and videos. These platforms not only serve as channels for sharing news and personal experiences but also for spreading rumors and shaping public opinions Ferrara (2020); Vosoughi et al. (2018). The inherent multimodality of social media content requires users to not only interpret individual modalities such as text or images but also to understand the interplay between them, pushing the boundaries of how machines comprehend human communication in online spaces.
Multimodal Large Language Models (MLLMs) have recently emerged as powerful tools for bridging the understanding of natural language and visual cues, showcasing their potential in a range of tasks ranging from image captioning to complex question answering Ramos et al. (2023); Liu et al. (2023c, b). Despite these advancements, the complexity of tasks such as understanding human emotions, memes, and verifying misinformation presents significant evaluation challenges to MLLMs. These tasks require not only combining signals extracted from both textual and visual domains, but also considering various social contexts upon making a decision regarding contextual appropriateness or correctness, which often require knowledge of cultural contexts and subjective interpretations Ruch (2010); Jacobi (2014). For instance, the task of explaining visual memes requires not only proficiency in image recognition and language generation, but also capability of understanding the underlying situation of the image on why it should be considered humorous. Given that large language models struggle at solving tasks requiring social knowledge Choi et al. (2023), we anticipate multimodal social tasks to prove an even harder challenge.
@@
Contributions. Our contributions are summarized as follows. First, we introduce MM-Soc, a novel benchmark to holistically evaluate MLLMs’ capability in tackling multimodal tasks derived from online social networks. Second, we perform a comprehensive evaluation and benchmark 10 representative open-source MLLMs on MM-Soc, comparing their performances with fine-tuned LLM baselines. Third, we conduct two case studies on MM-Soc for testing the effectiveness of two methods: self-improvement and explanation-augmented finetuning. We find that, while zero-shot MLLMs often fall short in achieving comparable performances compared to fine-tuned models, their performances can be improved via specific fine-tuning strategies. We aim to facilitate ongoing research and development in the field by releasing all of our code, data, and tools upon the acceptance of this work.
@@ 2 The MM-Soc Benchmark
Dataset
Domain
Modality
Size
PolitiFact
misinformation
news content, online posts,
images, user metadata
485
GossipCop
12,840
Hateful Memes
hate speech, OCR
images, embedded text
12,143
Memotion
sentiment, humor, OCR, offensiveness, sarcasm
images, embedded text
10,000
YouTube
tagging
images, text, channels
1,963,697
Table 1: Statistics of the MM-Soc benchmark.
@@
Motivated by this, we design MM-Soc, which spans both natural language understanding and generation tasks. These tasks are designed to test the models’ abilities to interact with user-generated content encountered online. The selection includes binary classification, multi-class classification, text extraction, and text generation tasks, aiming to cover a wide range of interactions MLLMs might encounter with online content. To ensure a comprehensive evaluation, we employ a variety of 10 tasks that mirror the complexity of real-world scenarios, from understanding online video contents, identifying misinformation to detecting hate speech in memes. The statistics of the dataset are in Table 1.
@@
Tagging. In digital content management, the ability to accurately predict appropriate tags for online content is particularly significant given their diverse and multimodal nature, which includes textual narratives, visual features, and cultural contexts. Effective tagging enhances content discoverability, facilitates content moderation, and significantly improves the user experience.
@@
To this end, we introduce YouTube2M, a novel dataset comprising 2 million YouTube videos, specifically curated to assess models’ proficiency in predicting tags from a predefined set in Table 7 based on video titles, descriptions, and visual content. We retrieved the URLs of all YouTube videos shared on Reddit over 12 years spanning from 2011 to 2022. Subsequently, we used YouTube Data API  [^1] https://developers.google.com/youtube/v3 to collect comprehensive metadata of the YouTube videos, including their titles, descriptions, channels, publish timestamps, restrictions, default languages, topic categories, and embeddability status. Additionally, we compiled extensive statistics for each video, covering aspects such as duration, and the number of comments, likes, and views they garnered. To ensure the quality and relevance of the dataset, we filtered the dataset and retained only videos with valid tags and thumbnail images, resulting in a dataset with 1,963,697 videos.
@@
Misinformation Detection. Misinformation detection represents a critical challenge as the proliferation of multimodal misinformation across online platforms can undermine trust in digital ecosystems and lead to real-world harm Yang et al. (2022, 2023); Jin et al. (2022); He et al. (2023); Ma et al. (2022). Here, we formulate misinformation detection as a binary classification problem and utilize the PolitiFact and GossipCop datasets Shu et al. (2020). The task aims at evaluating a model’s ability to accurately differentiate between true news and misinformation, leveraging both the textual content and the associated images of news articles.
@@ 3 Model Selection
Table 2: Performance comparison across all models on the tasks. Best and 2nd best performances among the MLLMs are highlighted in bold and underline, respectively. “ID” and “SCD” stand for the image description task and the social context description task, respectively. Note that instructblip-vicuna-7b fails to generate valid answers on the tagging task. A full comparison of all models on all metrics can be found in Appendix 8.
We consider 10 prominent open-source models spanning four different distinct architectures: LLaVA-v1.5 Liu et al. (2023b), BLIP2 Li et al. (2023b), InstructBLIP Dai et al. (2023), and LLaMA-Adapter-v2 Zhang et al. (2023b). Details on model parameter volumes are in Table 10. The models are selected to cover diverse model sizes. We apply our prompts (Table 6) to test the performances of MLLMs in a zero-shot setting. For tasks in which ground-truth texts are available as inputs, we compare MLLMs’ performances with five unimodal discriminative models in a full fine-tuning setting, including BERT Kenton and Toutanova (2019), RoBERTa-Base/Large Liu et al. (2019), DeBERTa He et al. (2020), and MiniLM Wang et al. (2020). These text-only models have shown competitive performances in text classification. Implementation details can be found in Appendix B.2.
@@ 4 Benchmark Results
Table 2 shows the overall performances across 10 tasks. Here, we use a unified score for each task to facilitate a high-level performance comparison across diverse tasks. For text classification and extraction tasks, we use the macro-F1 score as the aggregated measure. For text generation tasks including image description (ID) and social context description (SCD), we use ROUGE-L Lin (2004). The results for misinformation detection are averaged across PolitiFact and GossipCop, and the results for OCR are averaged across Memotion and Hateful Memes. The complete evaluation results can be found in Appendix 8.
@@ 5 Illustrative Uses of MM-Soc
PolitiFact GossipCop
[…]
llava-v1.5-13b 0.749 0.827 0.721 100.0 0.534 0.773 0.535 100.0
blip2-opt-2.7b 0.213 0.227 0.429 21.2 0.309 0.309 0.437 11.2
blip2-flan-t5-xl 0.419 0.721 0.500 100.0 0.514 0.819 0.534 100.0
blip2-flan-t5-xxl 0.545 0.548 0.634 100.0 0.200 0.215 0.481 100.0
llama-adapter-v2 0.550 0.553 0.613 87.5 0.556 0.673 0.581 83.6
roberta-base 0.894 0.923 0.894 100.0 0.812 0.879 0.824 100.0
roberta-large 0.846 0.885 0.825 100.0 0.820 0.858 0.820 100.0
MiniLM-v2 0.793 0.827 0.806 100.0 0.777 0.858 0.785 100.0
deberta-v3-large 0.952 0.962 0.952 100.0 0.817 0.895 0.792 100.0
random / 0.471 0.500 0.494 / 0.448 0.500 0.500 /
@@ 5.1 Can MLLMs Self-improve Its Answers?
The ability of MLLMs to self-improve – enhancing their answers iteratively without external supervised signals – can help generate increasingly consistent and robust answers, diminishing the need for human oversight. Using our benchmark, we investigate the self-improvement capabilities of MLLMs. The initial phase involves the model generating an answer for each question. Subsequent iterations, starting from the second round, require the model to produce new answers conditioned on the multimodal inputs and its prior responses. The iterative process is performed for six rounds. To quantitatively assess the evolution of answers across these iterations, we employed three established similarity metrics: BERTScore Zhang et al. (2019), sentence embeddings similarity Reimers and Gurevych (2019), and bigram similarity Kondrak (2005). These metrics enabled us to measure the consistency of answers from one round to the next, as well as their fidelity to the ground truth.
@@ 5.2 Does finetuning MLLMs Improve Overall Performance?
We employed two distinct strategies for fine-tuning. The first approach directly fine-tunes the model using the default input and output data, analogous to a standard fine-tuning setting. In the second approach, we leverage GPT-4(V) as a strong teacher to generate explanations after each ground truth answer for each sample. Along with the original input data, the GPT-generated explanations are augmented as additional training data.
Figure 5 shows the performances of fine-tuned LLaVA-7B and 13B models along with baselines; details can be found in Appendix B.3. With standard fine-tuning, we observe notable gains in detecting misinformation, offensiveness, and sentiment, but also drops in hate, humor, and sarcasm detection. Meanwhile, fine-tuning with explanations improved performance across a broader spectrum of tasks, e.g., increases of 18.2% in hate speech detection and 12.7% in sentiment analysis. Notably, text generation tasks such as image description and social context demonstrated greater gains.
Table 4 further reinforces the positive effects of finetuning with explanations for text generation tasks. Compared to the zero-shot baseline, both the 7B & 13B LLaVA models achieve higher ROUGE-2 scores on image description (40.5% for 7B and 30.5% for 13B). Similarly, for social context description, we observe improvements of 20.9% and 11.6% respectively. These improvements are accompanied by a reduction in response verbosity, highlighting the importance of explanations and rationales for improving multimodal text generation tasks. Interestingly, finetuning without explanations performs worse than the baseline, indicating that the standard finetuning approach may not be sufficient to learn the tasks in MM-Soc and signaling the need for refined finetuning strategies.
@@ 6 Related Works
Multimodal Large Language Models: Multimodal Large Language Models (MLLMs) have demonstrated exceptional natural language understanding and generation abilities by integrating visual information with textual inputs Awadalla et al. (2023); Yu et al. (2023); Liu et al. (2023a); Verma et al. (2023); Li et al. (2023c); Zeng et al. (2023). Models such as LLaVA Liu et al. (2023b, c), BLIP2 Li et al. (2023b), InstructBLIP Dai et al. (2023), and LLaMA-Adapter Zhang et al. (2023b); Gao et al. (2023) have showcased their superior performance in a range of applications. The success of MLLMs suggests their potential for widespread use in scenarios requiring not only factual analysis and comprehension but also subjective judgment and decision-making based on a nuanced understanding of social contexts and human perceptions. Our study reveals that current MLLMs still fall short in fully grasping and responding to complex social scenarios with the required depth of understanding and sensitivity.
Benchmarking Large Language Models: The evaluation of LLMs is crucial for uncovering their capabilities and identifying potential risks associated with their deployment in sensitive domains Wang et al. (2024, 2023); Liu et al. (2020); Zhang et al. (2023a); Zhao et al. (2023); Zong et al. (2023); Xiao et al. (2023); Chan et al. (2023). Benchmarking efforts across various domains such as legal Deroy et al. (2023), healthcare Jin et al. (2023a), finance Zhou et al. (2023), psychology Li et al. (2023a) have provided valuable insights into LLMs such as their reliability Shu et al. (2023), robustness Zhu et al. (2023b), and ethical implications Sun et al. (2023). Despite these efforts, there remains a notable gap in the development of comprehensive multimodal benchmarks for social domains. In this work, we create a holistic multimodal benchmark that captures the broad spectrum of social language and interactions.

### v2

@@ Abstract
Social media platforms are hubs for multimodal information exchange, encompassing text, images, and videos, making it challenging for machines to comprehend the information or emotions associated with interactions in online spaces. Multimodal Large Language Models (MLLMs) have emerged as a promising solution to these challenges, yet they struggle to accurately interpret human emotions and complex content such as misinformation. This paper introduces MM-Soc, a comprehensive benchmark designed to evaluate MLLMs’ understanding of multimodal social media content. MM-Soc compiles prominent multimodal datasets and incorporates a novel large-scale YouTube tagging dataset, targeting a range of tasks from misinformation detection, hate speech detection, and social context generation. Through our exhaustive evaluation on ten size-variants of four open-source MLLMs, we have identified significant performance disparities, highlighting the need for advancements in models’ social understanding capabilities. Our analysis reveals that, in a zero-shot setting, various types of MLLMs generally exhibit difficulties in handling social media tasks. However, MLLMs demonstrate performance improvements post fine-tuning, suggesting potential pathways for improvement. Our code and data are available at https://github.com/claws-lab/MMSoc.git.
@@ 1 Introduction
Social media platforms have become the epicenter of multimodal information exchange, blending various formats of content such as text, images, and videos. These platforms serve not only as channels for sharing news and personal experiences but also for spreading rumors and shaping public opinion Ferrara (2020); Vosoughi et al. (2018). The inherent multimodality of social media content requires users to not only interpret individual modalities such as text or images but also to understand the interplay between them, pushing the boundaries of how machines comprehend human communication in online spaces.
Multimodal Large Language Models (MLLMs) have recently emerged as powerful tools for bridging the understanding of natural language and visual cues, showcasing their potential in a range of tasks ranging from image captioning to complex question answering Ramos et al. (2023); Liu et al. (2023c, b). Despite these advancements, the complexity of tasks such as understanding human emotions, memes, and verifying misinformation presents significant evaluation challenges to MLLMs Chen and Shu (2023a, b). These tasks require not only combining signals extracted from both textual and visual domains, but also considering various social contexts upon making a decision regarding contextual appropriateness or correctness, which often require knowledge of cultural contexts and subjective interpretations Ruch (2010); Jacobi (2014). For instance, the task of explaining visual memes requires not only proficiency in image recognition and language generation, but also capability of understanding the underlying situation of the image on why it should be considered humorous. Given that large language models struggle at solving tasks requiring social knowledge Choi et al. (2023), we anticipate multimodal social tasks to prove an even harder challenge.
@@
Contributions. Our contributions are summarized as follows. First, we introduce MM-Soc, a novel benchmark to holistically evaluate MLLMs’ capability in tackling multimodal tasks derived from online social networks. Second, we perform a comprehensive evaluation and benchmark 10 representative open-source MLLMs on MM-Soc, comparing their performances with fine-tuned LLM baselines. Third, we conduct two case studies on MM-Soc for testing the effectiveness of two methods: self-improvement and explanation-augmented finetuning. We find that, while zero-shot MLLMs often fall short in achieving comparable performances compared to fine-tuned models, their performances can be improved via specific fine-tuning strategies.
@@ 2 The MM-Soc Benchmark
Motivated by this, we design MM-Soc, which spans both natural language understanding and generation tasks. These tasks are designed to test the models’ abilities to interact with user-generated content encountered online. The selection includes binary classification, multi-class classification, text extraction, and text generation tasks, aiming to cover a wide range of interactions MLLMs might encounter with online content. The detailed task selection process is in Appendix A. To ensure a comprehensive evaluation, we employ a variety of 10 tasks that mirror the complexity of real-world scenarios, ranging from understanding online video content to identifying misinformation and detecting hate speech in memes. The statistics of the dataset are in Table 5.
@@
Tagging. In digital content management, the ability to accurately predict appropriate tags for online content is particularly significant given their diverse and multimodal nature, which includes textual narratives, visual features, and cultural contexts. Effective tagging enhances content discoverability, facilitates content moderation, and significantly improves the user experience. To this end, we introduce YouTube2M, a novel dataset comprising 2 million YouTube videos shared on Reddit, specifically curated to assess models’ proficiency in predicting tags from a predefined set in Table 7 based on video titles, descriptions, and visual content. We provide a comprehensive analysis of YouTube2M in Appendix B.
YouTube2M distinguishes itself with two features: 1) Relevance to Online User Groups. The YouTube2M dataset features videos shared on Reddit. The selection of YouTube as the primary source is based on its expansive user base, with more than 2.5 billion monthly users Statista (2024) and the rich variety of its multimodal content. Reddit is among the top most popular social media and is characterized by its unique community structures called “subreddits”. Unlike general video collections on YouTube, YouTube2M reflects the choices of individuals within specific subreddit communities, aligning with their interests, humor, or preferences. This targeted selection process ensures our dataset is particularly relevant to distinct user groups. 2) Viral Potential. Reddit is renowned as a catalyst for virality. Videos shared on Reddit can rapidly gain significant attention and engage communities deeply through more discussions, comments, and votes within their respective subreddits. Notably, the presence of toxic, biased, or unverified content in online videos raises concerns over the propagation of misinformation, fostering distrust and hate speech online. Consequently, the accurate categorization and tagging of these videos become critical for content moderation.
@@
Dataset Construction. We retrieved the URLs of all YouTube videos shared on Reddit over 12 years spanning from 2011 to 2022. Subsequently, we used YouTube Data API  [^1] https://developers.google.com/youtube/v3 to collect comprehensive metadata of the YouTube videos, including their titles, descriptions, channels, publish timestamps, restrictions, default languages, topic categories, and embeddability status. Additionally, we compiled extensive statistics for each video, covering aspects such as duration, and the number of comments, likes, and views they garnered. To ensure the quality and relevance of the dataset, we filtered the dataset and retained only videos with valid tags and thumbnail images, resulting in a dataset with 1,963,697 videos.
@@
Misinformation Detection. Misinformation detection represents a critical challenge as the proliferation of multimodal misinformation across online platforms can undermine trust in digital ecosystems and lead to real-world harm Yang et al. (2022, 2023); Ma et al. (2022); Jin et al. (2022); He et al. (2023). Here, we formulate misinformation detection as a binary classification problem and utilize the PolitiFact and GossipCop datasets Shu et al. (2020). The task aims at evaluating a model’s ability to accurately differentiate between true news and misinformation, leveraging both the textual content and the associated images of news articles.
@@ 2.1 Model Selection
Table 1: Performance comparison across all models on the tasks. Best and 2nd best performances among the MLLMs are highlighted in bold and underline, respectively. “ID” and “SCD” stand for the image description task and the social context description task, respectively. Note that instructblip-vicuna-7b fails to generate valid answers on the tagging task. A full comparison of all models on all metrics can be found in Appendix D.2.
We consider 10 prominent open-source models spanning four different distinct architectures: LLaVA-v1.5 Liu et al. (2023b), BLIP2 Li et al. (2023b), InstructBLIP Dai et al. (2023), and LLaMA-Adapter-v2 Zhang et al. (2023b). Details on model parameter volumes are in Table 13. The models are selected to cover diverse model sizes. We apply our prompts (Table 6) to test the performances of MLLMs in a zero-shot setting. For tasks in which ground-truth texts are available as inputs, we compare MLLMs’ performances with five unimodal discriminative models in a full fine-tuning setting, including BERT Kenton and Toutanova (2019), RoBERTa-Base/Large Liu et al. (2019), DeBERTa He et al. (2020), and MiniLM Wang et al. (2020). These text-only models have shown competitive performances in text classification. Implementation details can be found in Appendix D.1.
@@ 3 Benchmark Results
Table 1 shows the overall performances across 10 tasks. Here, we use a unified score for each task to facilitate a high-level performance comparison across diverse tasks. For text classification and extraction tasks, we use the macro-F1 score as the aggregated measure. For text generation tasks including image description (ID) and social context description (SCD), we use ROUGE-L Lin (2004). The results for misinformation detection are averaged across PolitiFact and GossipCop, and the results for OCR are averaged across Memotion and Hateful Memes. The complete evaluation results can be found in Appendix D.2.
@@
Model Pre Rec F1 Jacc $\mathcal{L}_{\mathrm{Ham}}$downarrow
instructblip-vicuna-13b 0.044 0.230 0.050 0.032 0.429
instructblip-flan-t5-xl 0.045 0.326 0.057 0.036 0.500
instructblip-flan-t5-xxl 0.092 0.376 0.128 0.078 0.161
blip2-opt-2.7b 0.027 0.037 0.022 0.013 0.223
blip2-flan-t5-xl 0.196 0.191 0.157 0.112 0.092
blip2-flan-t5-xxl 0.176 0.350 0.188 0.122 0.085
llama-adapter-v2 0.028 0.029 0.021 0.012 0.137
llava-v1.5-7b 0.048 0.345 0.068 0.041 0.406
+ FT 0.162 0.373 0.209 0.148 0.063
+ FT w/ explanations 0.562 0.491 0.494 0.400 0.027
llava-v1.5-13b 0.052 0.361 0.071 0.043 0.342
+ FT 0.123 0.441 0.167 0.113 0.104
+ FT w/ explanations 0.533 0.473 0.474 0.387 0.027
Table 4: Results of tagging on the YouTube dataset. A $\downarrow$” next to the metric indicates that lower values represent better performances. instructblip-vicuna-7b fails to produce valid predictions in this context.
@@ 4.1 Can MLLMs Self-improve Its Answers?
The ability of MLLMs to self-improve – enhancing their answers iteratively without external supervision – can help generate increasingly consistent and robust answers, reducing the need for human oversight. Using our benchmark, we investigate the self-improvement capabilities of MLLMs. The initial phase involves the model generating an answer for each question. Subsequent iterations, starting from the second round, require the model to produce new answers conditioned on the multimodal inputs and its prior responses. The iterative process is performed for six rounds. To quantitatively assess the evolution of answers across these iterations, we employed three established similarity metrics: BERTScore Zhang et al. (2019), sentence embeddings similarity Reimers and Gurevych (2019), and bigram similarity Kondrak (2005). These metrics enabled us to measure the consistency of answers from one round to the next, as well as their fidelity to the ground truth.
@@ 4.2 Does finetuning MLLMs Improve Overall Performance?
We employed two distinct strategies for fine-tuning. The first approach directly fine-tunes the model using the default input and output data, analogous to a standard fine-tuning setting. In the second approach, we leverage GPT-4V as a strong teacher to generate explanations after each ground truth answer for each sample. Along with the original input data, the GPT-generated explanations are augmented as additional training data.
Figure 5 shows the performances of fine-tuned LLaVA-7B and 13B models along with baselines. With standard fine-tuning, we observe notable gains in detecting misinformation, offensiveness, and sentiment, but also drops in hate, humor, and sarcasm detection. Meanwhile, fine-tuning with explanations improved performance across a broader spectrum of tasks, e.g., increases of 18.2% in hate speech detection and 12.7% in sentiment analysis. Notably, text generation tasks such as image description and social context demonstrated greater gains. Table 3 further reinforces the positive effects of finetuning with explanations for text generation tasks. Compared to the zero-shot baseline, both the 7B & 13B LLaVA models achieve higher ROUGE-2 scores on image description (40.5% for 7B and 30.5% for 13B). Similarly, for social context description, we observe improvements of 20.9% and 11.6% respectively. These improvements are accompanied by a reduction in response verbosity, highlighting the importance of explanations and rationales for improving multimodal text generation tasks. Interestingly, finetuning without explanations performs worse than the baseline, indicating that the standard finetuning approach may not be sufficient to learn the tasks in MM-Soc and signaling the need for refined finetuning strategies.
@@ 5 Related Works
Multimodal Large Language Models: Multimodal Large Language Models (MLLMs) have demonstrated exceptional natural language understanding and generation abilities by integrating visual information with textual inputs Awadalla et al. (2023); Yu et al. (2023); Liu et al. (2023a); Verma et al. (2023); Li et al. (2023c); Zeng et al. (2023); Xie et al. (2024). Models such as LLaVA Liu et al. (2023b, c), BLIP2 Li et al. (2023b), InstructBLIP Dai et al. (2023), and LLaMA-Adapter Zhang et al. (2023b); Gao et al. (2023) have showcased their superior performance in a range of applications. The success of MLLMs suggests their potential for widespread use in scenarios requiring not only factual analysis and comprehension but also subjective judgment and decision-making based on a nuanced understanding of social contexts and human perceptions. Our study reveals that current MLLMs still fall short of fully grasping and responding to complex social scenarios with the required depth of understanding and sensitivity.
Benchmarking Large Language Models: The evaluation of LLMs is crucial for uncovering their capabilities and identifying potential risks associated with their deployment in sensitive domains Wang et al. (2024a, b, 2023); Liu et al. (2020); Zhang et al. (2023a); Zhao et al. (2023); Zong et al. (2023); Xiao et al. (2023); Chan et al. (2023). Benchmarking efforts across various domains such as legal Deroy et al. (2023), healthcare Jin et al. (2023), finance Zhou et al. (2023), psychology Li et al. (2023a) have provided valuable insights into LLMs such as their reliability Shu et al. (2023), robustness Zhu et al. (2023b), and ethical implications Sun et al. (2023). Despite these efforts, there remains a notable gap in the development of comprehensive multimodal benchmarks for social domains. In this work, we create a holistic multimodal benchmark that captures the broad spectrum of social language and interactions.